OpenStudy (anonymous):

Find the polar equation of the circle which passes through the pole and the two points whose polar coordinates are (d,0) and (2d,r/3)

OpenStudy (angela210793):

I don't remember them at all...my hs teacher never paid attention to these :( Sorry...

OpenStudy (eyust707):



OpenStudy (eyust707):

not exactly sure how to come up with an equation. that satisfies that... seems odd?

OpenStudy (eyust707):

you certinly could...

OpenStudy (anonymous):

(d,0) and (2d cosr/3, 2d sinr/3) \[x^2 + y^2 + 2gx + 2fy = 0\]

OpenStudy (anonymous):

As the circle passes from the origin the Constant is zero.

OpenStudy (anonymous):

x-Intercept and y intercept of the circle are -2g, -2f respectively. Since -2g= d, g = -d/2.

OpenStudy (anonymous):

Now get the points in the equation and you will get the value of f.

OpenStudy (anonymous):

It would have been neater and easier if r was pi.

OpenStudy (anonymous):

I'm sure @FoolForMath has a better way to do it.

OpenStudy (dumbcow):

you could also start with a general polar equation for a circle \[r = Acos(\theta)+Bsin(\theta)\] plug in the 2 given points to find A and B \[A = d\] \[B = \frac{d(2-\cos(r/3))}{\sin(r/3)}\] also i'm assuming r is just a constant and not the dependent variable "r" used in polar equations